10) (4 points) Prove or disapprove the function f:R → R such f(x) = 3x - 2 is one-to-one?
Answer #1

f!IR- IR. f(x) = 3x-2. X, G2 EIR let f(xi) = f(1/2) 3) 374-2= 382-2 7 34 = 342 24 =12 f is one-one. C Scanned with CamScanner